the house plans are 60 x 70 you decrease both length and width by the same amount, the area is 3696. so how much are the both decreased by? i need to know how to solve it.
Answer by ankor@dixie-net.com(22740)

the house plans are 60 x 70 you decrease both length and width by the same amount, the area is 3696.
so how much are the both decreased by? i need to know how to solve it.
:
Let x = amt decreased
:
(60 - x) * (70 - x) = 3696
FOIL
4200 - 60x - 70x + x^2 = 3696
Arrange as a quadratic equation
x^2 - 130x + 4200 - 3696 = 0
:
x^2 - 130x + 504 = 0
(x - 126)(x - 4) = 0
Two solutions
x = 126
and
x = 4, this is the solution that makes sense,
:
The new dimensions: 56 by 66
:
Check the area 56 * 66 = 3696
: